OpenStudy (anonymous):

If the vectors A and B represented in such a way that A+B=A-B, then the angle b/w A and B is ?

OpenStudy (anonymous):

Any idea @oksuz_ ?

OpenStudy (anonymous):

You can use Cosine's law for that...

OpenStudy (anonymous):

How ? Can u solve it please ?

OpenStudy (anonymous):

@AllTehMaffs Any idea ?

OpenStudy (anonymous):

For A+B A+B= A^2 + B^2 - 2*A*B*Cos(180-alpha)= A^2 + B^2 + 2*A*B*Cos(alpha) For A-B A-B = A^2 + B^2 - 2*A*B*Cos(alpha) if you make them equal each other.. 2*A*B*Cos(alpha) =- 2*A*B*Cos(alpha) 2AB vanishes both side Cos(alpha) = - Cos(alpha) 2Cos(alpha) = 0 this equation valid if and only if Cos(alpha)= (n*pi)/2 where n=1,3,5... so if we put n= 1 you get alpha= pi/2 that means A and B are perpendicular each other..

OpenStudy (anonymous):

Means 90 degree C @oksuz_ ?

OpenStudy (anonymous):

Sorry not C .. :P

OpenStudy (anonymous):

yes 90 degrees but i am not sure whether it is C or not :)

OpenStudy (anonymous):

No it is not ..mii mistake :P

OpenStudy (anonymous):

@AllTehMaffs right ans is 90 But Could n't figure out how ... Thanks @oksuz_ and @AllTehMaffs :)

OpenStudy (anonymous):

A+B= A^2 + B^2 - 2*A*B*Cos(180-alpha) . How Cos is (180 - alpha )? @oksuz_

OpenStudy (anonymous):

Cos(180-alpha)= Cos(180)*Cos(alpha) + Sin(180)*Sin(alpha) sin's terms vanishes and Cos(180)=-1

OpenStudy (anonymous):

ok .Same thing for A-B ?

OpenStudy (anonymous):

For A-B you get directly Cos(alpha)

OpenStudy (anonymous):

ok ty so much :)

OpenStudy (anonymous):

u are welcome :)